OpenStudy (anonymous):

what is the value of x in the equation 3^x =243 (1) 5 (2) 4 (3) 3 (4) 2

OpenStudy (whpalmer4):

3^2 = 3*3 = 3^3 = 3*3*3 = 3^4 = 3*3*3*3= 3^5 = 3*3*3*3*3=

OpenStudy (anonymous):

what

OpenStudy (whpalmer4):

Do the arithmetic. The answer to your problem should become apparent.

OpenStudy (anonymous):

It's 3^5

OpenStudy (anonymous):

( Therefore, it's (1) )